Ms. Hansen wants to install a 25 inch diagonal plasma TV screen into a cutout in her paneling. The screen in 20 inches wide. She cut a 20 inch by 15 inch rectangular opening in the paneling. Will the TV fit.
Answer by checkley71(8403)

THE ANSWER IS YES!
YOU HAVE A TRIANGLE THAT HAS AN HYPOTENUSE OF 25 INCHES & 1 SIDE OF 15 INCHES
25^2=20^2+15^2
625=400+225
625=625 THIS SAYS THE TV WILL FIT IF THE 25 INCHES IS AN OUTSIDE DIMENTION.